# Point-Spread-Function Engineering in MINFLUX: Optimality of Donut and Half-Moon Excitation Patterns

### v1

@@ B.2 Efficient Computation of the Generalized Inverse Fourier Transform
We see from Sections A.1 and A.4 that the inverse Fourier transform (IFT) is frequently required to compute the beam-propagation model and the partial derivatives found in the Fisher matrix. There are two notably aspects to consider in our specific case. First, the sizes of the input and output of the IFT can be different. Moreover, we would like to zoom in onto the object plane to refine PSF function (partial sampling). The ordinary IDFT is unfortunately unable to satisfy either aspect. In this section, we address the shortcomings of the ordinary IDFT by using the chirp Z-transform (CZT) and provide an efficient algorithm via the CZT. It has the same computational complexity as the FFT.
@@ C.2 Optimization for a Single Direction
To extend our approach to 2D, we first choose a direction defined by a unit vector $\mathbf{n}$. The directional gradient $D_{\mathbf{n}}g(\mathbf{r})\in\mathbb{C}$ of $g$ along a given direction $\mathbf{n}$ at $\mathbf{r}=\mathbf{0}$ is defined as
@@
Equality is achieved when $P(\mathbf{k})(\mathbf{k}\cdot\mathbf{n})=\mathrm{e}^{\mathrm{j}\alpha}$ for a constant $\alpha$ for all $\mathbf{k}\in\Lambda$. This yields that $P(\mathbf{k})=\mathrm{e}^{-\mathrm{j}\operatorname{Arg}(\mathbf{k}\cdot\mathbf{n})+\alpha}$ and corresponds to a half-moon pupil function as depicted in Fig. 6(a). This half-moon pupil function is characterized by a $\pi$ phase jump between the two half disks. It yields the optimal PSF that maximizes the gradient of the electric field in a single direction.
If one wants to minimize a single-directional gradient norm, the optimal pupil function corresponds to a half-moon PSF. This supports the result of our CRB optimization as we obtain 2 PSFs that maximize gradients along orthogonal directions to solve our 2D localization-estimation problem.
@@ C.3 Optimization Averaged Over All Directions
We see in Fig. 6(b) the results of these power iterations. They yield the phase ramp phase characteristics of the donut PSF. The integral over $\mathbf{n}\in O$ is discretized uniformly over the unit circle with $100$ terms. The pupil function $p\in\mathcal{H}$ is discretized over a uniform grid for $(k_{\text{x}},k_{\text{y}})$, as a $(2048\times 2048)$ px image. A total of $100$ power iterations is performed.
@@
If one wants to minimize directional gradient norm averaged along all directions, the optimal pupil function corresponds to a donut PSF. This supports the result of our CRB optimization as we obtain the donut PSF when only a single shifted PSF is allowed.
@@ D.3 Other Supporting Tables and Figures
In Fig. 10, we show the evolution of the loss and the run time of the optimization algorithm for all the experiments presented in the manuscript and the appendix.

### v2

@@ B.2 Efficient Computation of the Generalized Inverse Fourier Transform
We see from Sections A.1 and A.4 that the inverse Fourier transform (IFT) is frequently required to compute the beam-propagation model and the partial derivatives found in the Fisher matrix. There are two notable aspects to consider in our specific case. First, the sizes of the input and output of the IFT can be different. Moreover, we would like to zoom in onto the object plane to refine PSF function (partial sampling). The ordinary IDFT is unfortunately unable to satisfy either aspect. In this section, we address the shortcomings of the ordinary IDFT by using the chirp Z-transform (CZT) and provide an efficient algorithm via the CZT. It has the same computational complexity as the FFT.
@@ C.2 Optimization for a Single Direction
To extend our approach to 2D, we first choose a direction defined by a unit vector $\mathbf{n}$. The directional derivative $D_{\mathbf{n}}g(\mathbf{r})\in\mathbb{C}$ of $g$ along a given direction $\mathbf{n}$ at $\mathbf{r}=\mathbf{0}$ is defined as
@@
Equality is achieved when $P(\mathbf{k})(\mathbf{k}\cdot\mathbf{n})=\mathrm{e}^{\mathrm{j}\alpha}$ for a constant $\alpha$ for all $\mathbf{k}\in\Lambda$. This yields that $P(\mathbf{k})=\mathrm{e}^{-\mathrm{j}\operatorname{Arg}(\mathbf{k}\cdot\mathbf{n})+\alpha}$ and corresponds to a half-moon pupil function as depicted in Fig. 6(b). This half-moon pupil function is characterized by a $\pi$ phase jump between the two half disks. It yields the optimal PSF that maximizes the gradient of the electric field in a single direction.
If one wants to minimize a single-directional derivative norm, the optimal pupil function corresponds to a half-moon PSF. This supports the result of our CRB optimization as we obtain 2 PSFs that maximize gradients along orthogonal directions to solve our 2D localization-estimation problem.
@@ C.3 Optimization Averaged Over All Directions
We see in Fig. 6(d) the result of these power iterations. They yield the phase ramp phase characteristics of the donut PSF. The integral over $\mathbf{n}\in O$ is discretized uniformly over the unit circle with $100$ terms. The pupil function $p\in\mathcal{H}$ is discretized over a uniform grid for $(k_{\text{x}},k_{\text{y}})$, as a $(2048\times 2048)$ px image. A total of $100$ power iterations is performed.
@@
If one wants to minimize directional derivative norm averaged along all directions, the optimal pupil function corresponds to a donut PSF. This supports the result of our CRB optimization as we obtain the donut PSF when only a single shifted PSF is allowed.
@@ D.3 Other Supporting Tables and Figures
In Fig. 10, we show the evolution of the loss and the run time of the optimization algorithm for all the experiments presented in the manuscript and the Supplement.